# Now at ]Western Digital Research Center, Western Digital Corporation, San Jose, CA 95119, USA

### v1

@@ III Results
If the unconventional torque originated from the magnetic ordering of the CrPt3 we would expect the spin polarization to have a component along the magnetization direction. This would result in a purely conventional spin polarization in the device with current along the $[001]$ direction since the magnetization of the CrPt3 is perpendicular to the current direction, which is what we observed in our measurements. However, with current along the $[\overline{1}10]$ direction, the spin polarization should have an unconventional component along the current direction and collinear to the magnetization. This is not the case for our measurements where we only observe an unconventional spin polarization component with current along the $[\overline{1}11]$ and $[1\overline{1}1]$ directions. This alone suggests that the origin of the unconventional torque is not the magnetic ordering of the CrPt3. Additionally, since we observe an unconventional torque component even in disordered paramagnetic CrPt3, where the magnetic ordering is absent, the unconventional torque cannot come from the magnetic ordering.
Instead, it originates from the broken mirror symmetry due to the crystal structure of CrPt3. When current is applied along the $[1\overline{1}1]$ and $[\overline{1}11]$ directions it flows parallel to a low symmetry axis with a broken mirror symmetry [see Fig. 2(c)]. In contrast, when current is applied along the $[001]$ and $[\overline{1}10]$ directions it flows parallel to a high symmetry axis with the mirror symmetry preserved. The broken symmetry lifts the restriction that charge current, spin current, and spin polarization must be mutually perpendicular and allows unconventional spin polarization directions to exist [19]. Unconventional torques from a broken mirror symmetry due to the crystal structure have been observed in WTe2 [5], MnPd3 [20], and CoPt/CuPt [9], where an unconventional spin polarization is only generated with a current along a low symmetry axis. These results indicate that unconventional spin-orbit torques can be generated even in high symmetry materials as long as the current flow is along a low symmetry crystalline orientation.
@@ IV Conclusion
In summary, we have shown that an unconventional in-plane field-like torque is generated in both ordered (ferrimagnetic) and disordered (paramagnetic) CrPt3. Our spin-torque ferromagnetic resonance measurements reveal that only when current is applied parallel to axes where mirror symmetry is broken ($[1\overline{1}1]$ and $[\overline{1}11]$) is an unconventional torque generated. With current applied parallel to high symmetry axes ($[001]$ and $[\overline{1}10]$) only conventional torques are observed.
@@ V Acknowledgement
This work was supported as part of Quantum Materials for Energy Efficient Neuromorphic Computing (Q-MEEN-C), an Energy Frontier Research Center funded by the US Department of Energy, Office of Science, Basic Energy Sciences under Award No. DE-SC0019273.

### v2

@@ III Results
If the unconventional spin torque originated from the magnetic ordering of the CrPt3 the spin polarization is expected to have a component along the magnetization direction. This would result in a purely conventional spin polarization in the device with current along the $[001]$ direction since the magnetization of the CrPt3 is perpendicular to the current direction. With current along the $[\overline{1}10]$ direction, the spin polarization should have the largest unconventional spin-polarization component along the current direction and collinear to the magnetization. This is not the case for our measurements where we only observe an unconventional spin polarization component with current along the $[\overline{1}11]$ and $[1\overline{1}1]$ directions. This alone suggests that the origin of the unconventional torque is not the magnetic ordering of the CrPt3. Additionally, since we observe an unconventional torque component even in disordered paramagnetic CrPt3, where the magnetic ordering is absent, the unconventional torque cannot come from the magnetic ordering.
Unconventional spin-orbit torques arise when the system breaks the appropriate mirror plane symmetry, where the mirror plane is parallel to the electric field $\mathbb{E}$ and the out-of-plane vector $\mathbb{z}$. Breaking this symmetry enables spin currents with flow along $\mathbb{z}$ and spin polarization along $\mathbb{E}$ and $\mathbb{z}$. In the following, we denote spin currents by $Q_{ij}$ for $i,j\in[X,Y,Z]$, where $i$ is the spin flow direction and $j$ is the spin polarization direction. Assuming that the electric field points along $X$, the unconventional spin-orbit torques may arise from the spin currents $Q_{ZX}$ and $Q_{ZZ}$.
Figure 7: Spin-dependent Boltzmann calculations of the CrPt3/Cu/Py trilayer with quantum coherent boundary conditions. (a) In-plane charge current $j_{X}$ and spin-polarized current $Q_{XX}$ versus out-of-plane position $z$, where $\hat{\mathbb{M}}_{\text{Py}}=\hat{\mathbb{x}}$. (b)-(c) damping-like (cyan curves) and field-like (purple curves) torques versus in-plane magnetization direction $\phi$ for different $\Delta$ values, where the fitted curves are defined using Eq. 3. Unconventional spin-orbit torques refer to nonvanishing torques when $\hat{\mathbb{M}}_{\text{Py}}~{}||~{}\hat{\mathbb{z}}\times\mathbb{E}$ (here at $\phi=90\degree$), and are seen (as expected by symmetry arguments) when $\Delta\neq 0$. (d) Conventional (Y) and unconventional (X) torques, defined by Eqs. (4)-(5) as a function of $\Delta$. For CrPt3/Cu interfaces with broken mirror plane symmetry (i.e., $\Delta\neq 0$), both the conventional and unconventional field-like torques are larger than the damping-like torques.
In the samples measured, the CrPt3/Cu interface breaks the appropriate mirror plane symmetry while the Py layer, which receives the torque, does not. Thus, the microscopic mechanism for unconventional torques likely involve interlayer scattering. If the CrPt3/Cu layer or the CrPt3/Cu interface generates the spin current $Q_{ZX}$, a damping-like torque $\bm{\tau}\propto\mathbb{M}_{\text{Py}}\times(\mathbb{x}\times\mathbb{M}_{\text{Py}})$ results in Py via the spin transfer mechanism. An unconventional field-like torque, given by $\bm{\tau}\propto\mathbb{M}_{\text{Py}}\times\mathbb{x}$, results if the incident spin current’s polarization, initially along $\mathbb{x}$, is rotated about the magnetization, such that the polarization now has a component along $\mathbb{M}_{\text{Py}}\times\mathbb{x}$. The resulting spin transfer torque from the new component is $\bm{\tau}\propto\mathbb{M}_{\text{Py}}\times[(\mathbb{M}_{\text{Py}}\times\mathbb{x})\times\mathbb{M}_{\text{Py}}]=\mathbb{M}_{\text{Py}}\times\mathbb{x}$, which is field-like.
The rotation of the spin current polarization, which results in field-like torques, can occur in two ways. First, if the Cu/Py interface has a nonvanishing imaginary part of the spin mixing conductance, the incident spin current’s polarization is rotated about the magnetization upon reflection. However, for this mechanism to explain a ratio of field-like to damping-like torque greater than one, as seen in the experimental data, the ratio of imaginary to real part of the spin mixing conductance would also be greater than one, which is extremely unlikely at metallic interfaces. The other mechanism, called an indirect nonlocal spin-orbit torque [22], results from a combination of interlayer scattering and spin-to-spin conversion at the interface.
In the indirect, nonlocal mechanism, the in-plane, spin-polarized current in Py “leaks” through the spacer layer into the CrPt3 layer. This leakage is the well-known mechanism behind the current-in-plane giant magnetoresistance and is readily modeled by the spin-dependent Boltzmann equation, as seen in Fig. 7(a). The leaked, in-plane, spin current is polarized along $\mathbb{M}_{\text{Py}}$. At the CrPt3/Cu interface, spin-orbit scattering results in spin-to-spin conversion, where the leaked, in-plane spin current is converted into an out-of-plane spin current with spin polarization $\mathbb{M}_{\text{Py}}\times\mathbb{f}$, where the vector $\mathbb{f}$ is a function of the tensor elements that describe spin-to-spin conversion at the interface. The spin-to-spin conversion can result from spin-orbit precession [23, 24] at the CrPt3/Cu interface or from spin swapping [25] in the CrPt3 layer. The emitted spin current then flows through Cu into Py and exerts a spin-transfer torque given by $\bm{\tau}\propto\mathbb{M}_{\text{Py}}\times[(\mathbb{M}_{\text{Py}}\times\mathbb{f})\times\mathbb{M}_{\text{Py}}]=\mathbb{M}_{\text{Py}}\times\mathbb{f}$. The vector $\mathbb{f}$ has a nonvanishing $x$ and $y$ components in general (these components are quantified in free-electron models in [22]), thus resulting in both conventional and unconventional field-like torques.
We use a multilayer, spin-dependent Boltzmann formalism with quantum coherent boundary conditions to demonstrate that indirect, nonlocal spin-orbit torques plausibly explain the experimental results. Solving the multilayer Boltzmann equation yields the nonequilibrium charge distribution function $g_{c}(z,\mathbb{k})$ and nonequilibrium spin distribution function $g_{s,l}(z,\mathbb{k})$ for $l\in[x,y,z]$, which are used to compute the spin/charge currents and spin torques. We model the CrPt3/Cu/Py trilayer assuming a free electron dispersion relation within each layer and use the spin-dependent relaxation time approximation to construct the collision integrals. All relevant details of the theoretical formalism are provided in [22], though here we have added diffusive scattering at the interfaces to model disorder.
The interfacial scattering potential is given by:
$V(\mathbb{r},\mathbb{k})$ $=\frac{\hbar^{2}k_{\text{F}}}{2m}\delta(z)\Big{(}\mu_{0}+\mu_{\text{ex}}\bm{\sigma}\cdot\hat{\mathbb{M}}_{\text{Py}}$
$+(\mu_{\text{R}}+\Delta/2)\sigma_{x}k_{y}-(\mu_{\text{R}}-\Delta/2)\sigma_{y}k_{x}\Big{)}$ (2)
where $k_{\text{F}}$ is the Fermi momentum, $m$ is the effective electron mass, and $\mu_{0}$, $\mu_{\text{ex}}$, and $\mu_{\text{R}}$ are dimensionless parameters specifying the strength of the spin-independent potential barrier, exchange interaction, and Rashba spin-orbit coupling respectively. The dimensionless parameter $\Delta$ specifies the degree of mirror-plane symmetry-breaking at the interface, where $\Delta=0$ corresponds to an interface with the mirror plane symmetry intact. Aside from spin-flip scattering in the bulk layers (which causes spin diffusion but not spin current generation), spin-orbit coupling is turned off everywhere in the system except the CrPt3/Cu interface. Thus, we isolate nonlocal spin-orbit torques due to interlayer scattering, removing all contributions from bulk spin current generation (like the spin Hall effect, spin anomalous Hall effect, and the magnetic spin Hall effect) and inverse spin galvanic effects at the Cu/Py interface.
Fig. 7(a) shows the in-plane charge current $j_{X}$ and spin-polarized current $Q_{XX}$ as a function of out-of-plane position $z$, where $\hat{\mathbb{M}}_{\text{Py}}=\hat{\mathbb{x}}$. The spin-polarized current “leaks” into the CrPt3 layer, even changing signs due to spin-dependent scattering at the interfaces. Figs. 7(b)-(c) show the damping-like torques $\tau_{d}$ (cyan curves) and field-like torques $\tau_{f}$ (purple curves) as a function of in-plane $\hat{\mathbb{M}}_{\text{Py}}$ direction ($\phi$). When the system possesses the mirror plane symmetry, i.e., $\Delta=0$, the total spin torque must vanish when $\hat{\mathbb{M}}_{\text{Py}}~{}||~{}\hat{\mathbb{z}}\times\mathbb{E}$ (here at $\phi=90\degree$). Since the magnetization is swept in-plane, the damping-like and field-like torques are the in-plane and out-of-plane torque components respectively, and are fitted to the following form:
$\tau^{\text{fit}}_{d/f}$ $=A_{d/f}\cos(\phi+b_{d/f}).$ (3)
To define the conventional (Y) and unconventional (X) torques, we use these relations:
$\tau^{Y}_{d/f}$ $=A_{d/f}\cos(b_{d/f})$ (4)
$\tau^{X}_{d/f}$ $=A_{d/f}\sin(b_{d/f}).$ (5)
Fig. 7(d) shows the conventional and unconventional torques versus symmetry-breaking parameter $\Delta$. Both the conventional and unconventional field-like torques are greater than the damping-like torques, showing reasonable agreement with the experimental results. While the field-like torque remains dominant under small changes in the parameters of the model, further research is needed to ascertain whether this property is generic for the torque generated by the indirect mechanism.
@@ IV Conclusion
In summary, we have shown that an unconventional in-plane field-like torque is generated in both ordered (ferrimagnetic) and disordered (paramagnetic) CrPt3. Our spin-torque ferromagnetic resonance and second harmonic Hall measurements reveal that we only observe an unconventional spin-torque component when current is applied along the $[1\overline{1}1]$ and $[\overline{1}11]$ crystal directions. When current is applied along the $[001]$ and $[\overline{1}10]$ crystal directions only conventional torques are observed. Our calculations reveal that symmetry breaking at the CrPt3/Cu interface combined with an indirect non-local spin-orbit torque leads to unconventional spin-torques.
@@ V Acknowledgement
This work was supported as part of Quantum Materials for Energy Efficient Neuromorphic Computing (Q-MEEN-C), an Energy Frontier Research Center funded by the US Department of Energy, Office of Science, Basic Energy Sciences under Award No. DE-SC0019273. KDB acknowledges support from the National Science Foundation under Grants No. DMR-1916275 and DMR-2324203. VA acknowledges support from the National Science Foundation under grants DMR-2105219 and ECCS-2236159.